# Distribution-informed Online Conformal Prediction

### v1

@@ A.1 Discussions on Online Mirror Descent
We remark that if we can obtain the estimated distribution function $\hat{F}_{t}$ in an online fashion, a direct way to combine it with eq.˜3 is to leverage online mirror descent [orabona2019modern, wang2025mirror]:
@@ A.2 Discussions on ACI
Adaptive Conformal Inference (ACI) [aci_gibbs2021adaptive] is another widely used online CP update besides eq.˜3. Let $\hat{Q}_{t}(\cdot)$ be the fitted quantiles of the non-conformity scores in calibration set $\mathcal{D}_{\mathrm{cal}}$:
@@ A.3 Discussions on the "Same-sign" Assumption
as long as $F_{t+1}$ is $L$-Lipschitz continuous and $\lambda_{t+1}>0$ is small.
@@ B.1 Proof of Proposition 1
Proposition 1. Assume that $\hat{F}_{t+1}(\hat{q}_{t+1})-(1-\alpha)$ and $F_{t+1}(\hat{q}_{t+1})-(1-\alpha)$ have the same sign, and $F_{t+1}$ is $L$-Lipschitz continuous. With a suitably small $\lambda>0$, we have:
@@
Denote $\epsilon_{t+1}=\sup_{q}|F_{t+1}(q)-\hat{F}_{t+1}(q)|$, $\delta_{t+1}=|\hat{F}_{t+1}(\hat{q}_{t+1})-(1-\alpha)|\leq 1-\alpha$, then
@@
Hence, even if the same-sign assumption in Proposition˜1 does not hold, the "instantaneous" performance of the refinement step is bounded by $\lambda_{t+1}$ and $\epsilon_{t+1}$.
@@ B.2 Regret Guarantee with Constant Learning Rates
We begin by presenting the Bregman Proximal Inequality [chen1993convergence]:
@@ Appendix C Implement of COP with Kernel-based CDF
Distribution
Drift
ECDF 90.6 7.07 6.89 90.0 7.09 6.97 90.9 7.30 6.99
@@
Amazon
Stock
ECDF 89.6 39.86 27.91 89.5 17.09 12.90 89.6 17.21 12.23
@@
Google
Stock
ECDF 89.7 49.72 42.09 89.6 19.87 17.04 89.3 30.25 28.24
@@
Electricity
Demand
ECDF 90.1 0.385 0.376 90.0 0.117 0.098 89.8 0.069 0.052
@@ Appendix D Learning Rate Selection in the Experiments
where the hyperparameter $\epsilon=0.1$ follows [angelopoulos2024online]. For conformal PID, ECI and COP:
@@
where $s_{t}$ is the non-conformity score at time $t$ and the window length $w=100$ follows [pid_angelopoulos2024conformal].
@@ Appendix E More Experimental results in simulation datasets
Variance
Changepoint
ACI 91.0 $\infty$ 11.06 91.0 $\infty$ 10.74 91.0 $\infty$ 10.90
@@
Extreme
Drfit
ACI 90.4 $\infty$ 79.01 89.7 $\infty$ 61.85 90.3 $\infty$ 70.06

### v2

@@ A.1 Discussions on Online Mirror Descent
We remark that if we can obtain the estimated distribution function $\hat{F}_{t}$ in an online fashion, a direct way to combine it with eq.˜3 is to leverage online mirror descent (Orabona, 2019; Wang et al., 2025):
@@ A.2 Discussions on ACI
Adaptive Conformal Inference (ACI) (Gibbs and Candès, 2021) is another widely used online CP update besides eq.˜3. Let $\hat{Q}_{t}(\cdot)$ be the fitted quantiles of the non-conformity scores in calibration set $\mathcal{D}_{\mathrm{cal}}$:
@@ A.3 Discussions on the "Same-sign" Assumption
as long as $F_{t+1}$ is $L$-Lipschitz and $\lambda_{t+1}>0$ is small.
@@ B.1 Proof of Proposition 1
Proposition 1. Assume that $\hat{F}_{t+1}(\hat{q}_{t+1})-(1-\alpha)$ and $F_{t+1}(\hat{q}_{t+1})-(1-\alpha)$ have the same sign, and $F_{t+1}$ is $L$-Lipschitz. With a suitably small $\lambda>0$, we have:
@@
Denote $\epsilon_{t+1}=\sup_{q}|F_{t+1}(q)-\hat{F}_{t+1}(q)|$$\delta_{t+1}=|\hat{F}_{t+1}(\hat{q}_{t+1})-(1-\alpha)|\leq 1-\alpha$, then
@@
Hence, even if the same-sign assumption in Proposition˜1 does not hold, the "instantaneous" performance of the refinement step is bounded by $\lambda_{t+1}$ and $\epsilon_{t+1}$. ∎
@@ B.2 Regret Guarantee with Constant Learning Rates
We begin by presenting the Bregman Proximal Inequality (Chen and Teboulle, 1993):
@@ Appendix C Implement of COP with Kernel-based CDF
Distribution Drift ECDF 90.6 7.07 6.89 90.0 7.09 6.97 90.9 7.30 6.99
@@
Amazon Stock ECDF 89.6 39.86 27.91 89.5 17.09 12.90 89.6 17.21 12.23
@@
Google Stock ECDF 89.7 49.72 42.09 89.6 19.87 17.04 89.3 30.25 28.24
@@
Electricity Demand ECDF 90.1 0.385 0.376 90.0 0.117 0.098 89.8 0.069 0.052
@@ Appendix D Learning Rate Selection in the Experiments
where the hyperparameter $\epsilon=0.1$ follows Angelopoulos et al. (2024). For conformal PID, ECI and COP:
@@
where $s_{t}$ is the non-conformity score at time $t$ and the window length $w=100$ follows Angelopoulos et al. (2023b).
@@ Appendix E More Experimental results in simulation datasets
Variance Changepoint ACI 91.0 $\infty$ 11.06 91.0 $\infty$ 10.74 91.0 $\infty$ 10.90
@@
Extreme Drfit ACI 90.4 $\infty$ 79.01 89.7 $\infty$ 61.85 90.3 $\infty$ 70.06